Jim is timed 16 times to measure his performance in minutes of a certain task. The average of these 16 observations came out to be 8.5 minutes with a standard deviation of 1.25 minutes. Obtain a 90% confidence interval for the actual mean performance of Jack's work.
Answer #1

Solution :

Given that,

t\alpha /2,df = 1.753

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.753 * (1.25 / \sqrt 16)

Margin of error = E = 0.55

The 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

8.5 - 0.55 < \mu < 8.5 + 0.55

7.95 < \mu < 9.05

(7.95 , 9.05)